In a 51-inch television,the length of the screen’s height is 25 inches ,what is the width?round your answer to the nearest tenth of an inch


1
Expert's answer
2021-10-24T18:06:52-0400

Here, TV screen sizes are specified according to their diagonal measurements, so if it's 25 inches tall,


Let w= width and h= height


Then

"w^2+25^2=51^2\\w^2=51^2-25^2\\w^2=2601-625\\w^2=1976\\w=\sqrt{1976}\\w=44.45\space inches"